Please return to NE43-366

department of electrical engineering and computer science

massachusetts institute of technology

Cambridge, Massachusetts 02139

Application for Approval of Minor in Area II (for courses not taken at MIT)

Please print legibly and in ink. Name of Applicant: Address for reply: Email address: Date: Title or Subject of Minor:

Non-MIT Course No. and title Where Taken No. of hours

Grade

1st Course name (Course number)

School No. of hours

Grade

2nd Course name (Course number)

School No. of hours

Grade

3rd Course name (Course number)

School No. of hours

Grade

Student’s signature Counselor’s approval Area Chairperson’s approval (initials)